OpenStudy (anonymous):

The table below shows some values of f(x) and g(x) for different values of x. x f(x) = 2x + 24 g(x) = 5(0.5)x -2 20 ? -1 ? ? 0 ? 5 1 ? 2.5 2 28 ? Complete the chart and determine the solution of the equation f(x) = g(x)? x = -2 x = -1 x = 1 x = 20

OpenStudy (anonymous):

@undeadknight26 @phi @ganeshie8 @sel95

undeadknight26 (undeadknight26):

Can u take a snap shot of the problem because its a but hard to figure out like this.

OpenStudy (anonymous):

ok this should work

undeadknight26 (undeadknight26):

For this we need to plug in the numbers for x: f(x) = 2x + 24 f(-1) = 2 x -1 + 24.

OpenStudy (anonymous):

Question really quick what does it mean when it says complete the chart to "solve" the equation???

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

@agent0smith @austinL @e.mccormick @RadEn @robtobey @tHe_FiZiCx99 @Taylor<3sRin help this person please XD

OpenStudy (the_fizicx99):

You can either plot them in a graphing calculator or plug in each x value and see which gives you a true statement.

OpenStudy (anonymous):

I just don't understand what im supposed to solve for :(

OpenStudy (the_fizicx99):

Wait, is that being raised to a power? g(x) = 5(0.5)^x?

OpenStudy (anonymous):

yes the power of x

OpenStudy (the_fizicx99):

f(x) = 2x + 24 use x = 2 and plug in 2 for every x, f(2) = 2(2) + 24 f(2) = 4 + 24 = 28 f(2) = 28 g(x) = 5(0.5)^x g(2) = 5(0.5)^2 g(2) = 5(0.25) g(2) = 1.25

OpenStudy (the_fizicx99):

Try the others on your own. and see which gives you a true statement, like 7 = 7 <-- true

OpenStudy (anonymous):

ok i know how to do that but i just dont know what i am trying to solve for?

OpenStudy (the_fizicx99):

You're trying to solve for each of those x values. Which ever gives you a true statement is the correct one.

OpenStudy (anonymous):

so would it be x = -1

OpenStudy (the_fizicx99):

no try again, it has to be equal like 20 = 20 or 16 = 16 They both have to have the same answer

OpenStudy (anonymous):

wait i think im doing it wrong could you explain it real quick?

OpenStudy (the_fizicx99):

explain what

OpenStudy (the_fizicx99):

f(-2) = 2(-2) + 24 f(-2) = 20 g(-2) = 5*4 g(-2) = 20 Its x = -2 because THEY BOTH EQUAL 20

OpenStudy (anonymous):

oh so to find that they were true they had to both equal the same wow im dumb... THank you so much for explaining!

OpenStudy (the_fizicx99):

np